OpenStudy (anonymous):

cosine x = 3.5/4.6 solve for x

OpenStudy (anonymous):

Take inverse cosine of both sides to get \(x = \cos^{-1}\left(\dfrac{3.5}{4.6}\right)\) and then use your calculator to approximate this answer. Does this make sense? :-)

OpenStudy (anonymous):

yeah i get it, any chance you can give me the flat out answer tho i lost my calculator haha

OpenStudy (anonymous):

i got 40.459

OpenStudy (anonymous):

ty(: